# AssumptionBox linecolor=blue, outerlinewidth=2pt, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=10pt, innerleftmargin=10pt, backgroundcolor=gray!20

### v1

@@ 2.2. Fairness quantification
3. Profile Generation Framework
This section outlines a structured approach for creating meaningful user profiles based on their music listening histories and watched movies. The goal is to refine the relevance of prompts used with directly-applied language models (LLMs) within recommendation systems. This work utilizes two datasets, MovieLens and LastFM, in the movie and music domains, where each user has engaged with a range of 100 to thousands of items. Incorporating all these activities into a single prompt is impractical due to prompt token length limitations and the potential introduction of noise. Therefore, the key insights behind our approach proposed in this work are to segment the prompt into different sections that capture both historical preferences and recent activity. We believe this strategy is particularly relevant for the sequential recommendation tasks employed in this study, where a user next activity is directly influenced by their previous interactions.
3.1. Formulation of User Profiles
A user profile is formulated by integrating three distinct components: the demographic prompt, the passion profile, and the recent interaction profile. Each component targets specific aspects of user data and interactions, tailored for use with RecLLMs.
Definition 3.1 (User Profile Components).
The combined user profile, denoted as $\mathcal{P}$, is a composite prompt defined as:
(4) $\mathcal{P}=\mathcal{D}+\mathcal{P}_{p}+\mathcal{P}_{i}$
  • $\mathcal{D}$ represents the demographic prompt. It incorporates user demographics and counterfactual user demographics to explore fairness scenarios, where the latter is introduced in this work.
    Example of Demographic Prompt:
    The user is a Female.” or “The user is Young.”
  • $\mathcal{P}_{p}$ denotes the passion profile. It aims to capture the long-term preferences and tastes of the target user, as derived from their historical interactions with the system. Given that users in recommendation datasets often explore and interact with numerous items, we employ sampling strategies to better reflect their tastes (cf. Section 3.1.1).
    Example of Passion Profile.
    The user has listened to the following songs in the past, organized as (Song - Artist): –: “Flume” by Bon Iver –: “Timothy Hay” by Mewithoutyou –: “The Fox, The Crow, And The Cookie” by Mewithoutyou –: “The Angel Of Death Came To David’S Room” by Mewithoutyou This selection reflects the user’s musical preferences.
    Note that the system allows the number of items included in the passion profile to be controlled automatically via a variable. In this work, we have set this number to a fixed value of 10.
  • $\mathcal{P}_{i}$ signifies the interaction profile. It captures the most recent and relevant user-system interactions. This profile provides a snapshot of the current and ongoing interests of the user, which is crucial for understanding their immediate subsequent interactions in sequential or temporally-oriented recommendation scenarios.
    Example of Recent Interaction Profile (ICL-1).
    Given the user has recently listened to the following songs in order: –: “Points Of Authority” by Linkin Park –: “Runaway” by Linkin Park –: “In The End” by Linkin Park –: “A Place For My Head” by Linkin Park –: “Senang” by Kane You should recommend: “Aftiteling: Hold On To The World” by Kane. What would be the top-10 suitable next recommendations after “Aftiteling: Hold On To The World” by Kane?
    Please note that the above describes a few-shot scenario with one example denoted as ICL-1. In this work, we test and try three different cases: zero-shot (00-shot), few-shot with one and two examples (ICL-$1$ and ICL-$2$). Please note that here we also define a context window to determine the number of examples to take from the user’s final interaction in the training, but use a fixed number = 5, in this work.
3.1.1. Passion Profile Sampling Strategies
The passion profile $\mathcal{P}_{p}$ is constructed using one of the following sampling strategies, which balance the frequency and recency of interactions, made by users with items:
  • Random Sampling: Uniform selection of tracks or movies to provide a stochastic view of user preferences.
  • Frequent Sampling: Prioritization of tracks/movies based on their frequency of playback, and rating provided emphasizing the main preferences of the user.
  • Recent-Frequent Sampling: This hybrid approach combines the recency and frequency of track interactions using a weighted score formula. Specifically, the score for each track is calculated as:
    (5) $=\omega\times\textbf{recency\_score}+(1-\omega)\times\textbf{count\_score}$_score
    where recency_score is a measure of how recent the user interactions with the track were, normalized to the range [0,1]—with more recent interactions scoring higher. count reflects the frequency of interactions, normalized similarly to the range [0,1]—with more frequent interactions scoring higher. $\omega$ is a configurable parameter (where $0\leq\omega\leq 1$) that balances the importance of recency versus frequency, aimed to fit the specific dynamics of the user-item interaction patterns or the recommendation strategy. In our experiments, we set $\omega=0.65$, thereby giving more importance to recent items.
The user profile is then created by combining the three prompts mentioned above and concatenating them to create a reasonably sized and informative user profile to guide recommendations.
3.2. Applications and Implications
The structured user profiles are utilized to assess the fairness and accuracy of recommendations in systems employing LLMs. By systematically varying demographic and interaction components, researchers can identify and mitigate potential biases inherent in algorithmic recommendations. These profiles thus serve both diagnostic and corrective functions in the ongoing effort to enhance AI-driven personalization in media consumption.
4. Evaluation methodology
Our framework is designed to rigorously assess fairness in Recommender Systems powered by Large Language Models (RecLLMs). We evaluate the models as follows:
  • On the consumer side, we assess fairness notions involving the sensitive ranker (NSD and SCSD), using a “benefit-driven” approach (cf. Section 2.1.3) with the neutral recommender ($\mathcal{R}_{N}$) serving as the reference.
  • Furthermore on the consumer side, we evaluate the Intrinsic Fairness (IF) notion, using a “uniform distribution” and “proportional distribution” as the target to measure fairness across different sensitive groups (cf. Section 2.1.3).
Table 2. Summary of Fairness Metrics for RecLLM Evaluation
def.
ICL $\Delta\mathcal{B}_{{a_{1}}}$ $\Delta\mathcal{B}_{{a_{2}}}$ $\delta_{\text{gender}}$ SSND SSNR $\Delta\mathcal{B}_{{a_{1}}}$ $\delta\mathcal{B}_{{a_{2}}}$ $\delta_{\text{gender}}$ SSND SSNR
Gender NSD 00-shot/rand. -0.0074 -0.019 0.0116 (p=0.730) 0.6667 0.1454 0.0 -0.0857 0.0857 (p=0.679) 6.0 0.8893
ICL-1/rand. -0.0222 -0.0476 0.0254 (p=0.464) 1.0 0.1527 -0.0333 -0.1571 0.1238 (p=0.389) 6.0 0.6373
ICL-2/rand. -0.037 -0.019 -0.018 (p=0.386) 0.3333 0.0921 -0.1556 -0.0429 -0.1127 (p=0.303) 4.0 0.4921
00-shot/freq. -0.0 -0.019 0.019 (p=0.609) 1.0 0.1599 0.0222 0.0143 0.0079 (p=0.962) 6.0 0.7276
ICL-1/freq. 0.0148 -0.0095 0.0243 (p=0.368) 1.0 0.1203 0.0889 -0.0143 0.1032 (p=0.301) 4.0 0.4537
ICL-2/freq. 0.0148 -0.019 0.0339 (p=0.425) 2.0 0.1908 0.0333 -0.0857 0.1190 (p=0.373) 5.0 0.5944
00-shot/rec-freq. -0.0148 0.0095 -0.0243 (p=0.587) 1.3333 0.1943 -0.0111 0.0 -0.0111 (p=0.946) 5.0 0.7218
ICL-1/rec-freq. 0.0444 0.0 0.0444 (p=0.241) 1.0 0.1682 0.$1$0.0143 0.1254 (p=0.411) 5.0 0.6514
ICL-2/rec-freq. -0.0074 -0.0 -0.0074 (p=0.837) 1.0 0.1598 0.0333 0.1286 -0.0952 (p=0.578) 6.0 0.7802
NCSD 00-shot/rand. 0.0286 -0.0 0.0286 (p=0.605) 1.6667 0.2363 -0.0714 0.0778 -0.1492 (p=0.544) 6.0 1.0772
ICL-1/rand. -0.0762 -0.0148 -0.0614 (p=0.169) 1.3333 0.1956 -0.2143 0.0778 -0.2921 (p=0.146) 6.0 0.8514
ICL-2/rand. -0.019 -0.0074 -0.0116 (p=0.738) 1.3333 0.157 -0.0571 -0.0778 0.0206 (p=0.855) 3.0 0.4917
00-shot/freq. -0.019 -0.0074 -0.0116 (p=0.793) 1.0 0.1933 0.0714 0.0 0.0714 (p=0.739) 6.0 0.9425
ICL-1/freq. -0.019 -0.0222 0.0032 (p=0.948) 1.6667 0.2152 0.0 -0.1333 0.1333 (p=0.415) 6.0 0.7198
ICL-2/freq. -0.0286 -0.0519 0.0233 (p=0.575) 1.3333 0.1831 -0.0857 -0.2 0.1143 (p=0.509) 6.0 0.7823
00-shot/rec-freq. 0.0381 -0.0593 0.0974 (p=0.062) 1.6667 0.2277 0.1286 -0.1556 0.2841 (p=0.127) 6.0 0.8275
ICL-1/rec-freq. 0.0095 -0.0074 0.0169 (p=0.749) 2.0 0.2418 0.0429 -0.0111 0.054 (p=0.786) 6.0 0.893
ICL-2/rec-freq. -0.019 0.0148 -0.0339 (p=0.343) 1.0 0.1516 0.1857 0.0778 0.1079 (p=0.553) 5.0 0.784
$\Delta\mathcal{B}_{{b_{1}}}$ $\Delta\mathcal{B}_{{b_{2}}}$ $\delta_{\text{age-gr.}}$ SSND SSNR $\Delta\mathcal{B}_{{b_{1}}}$ $\delta\mathcal{B}_{{b_{2}}}$ $\delta_{\text{age-gr.}}$ SSND SSNR
Age-group NSD 00-shot/rand. -0.0183 0.0 -0.0183 (p=0.814) 1.0 0.1576 -0.0685 -0.2857 0.2172 (p=0.666) 6.0 0.9526
ICL-$1$/rand. -0.0046 0.0476 -0.0522 (p=0.573) 0.6667 0.1673 -0.0068 0.7857 -0.7926 (p=0.137) 6.0 0.8694
ICL-$2$/rand. -0.0183 0.0476 -0.0659 (p=0.226) 1.0 0.1194 0.089 -0.2857 0.3748 (p=0.094) 4.0 0.5039
00-shot/freq. -0.0228 0.0 -0.0228 (p=0.228) 1.0 0.1126 -0.0959 -0.1429 0.047 (p=0.784) 5.0 0.575
ICL-$1$/freq. -0.0091 0.0 -0.0091 (p=0.483) 0.6667 0.0778 0.0137 0.0 0.0137 (p=0.754) 3.0 0.2615
ICL-$2$/freq. -0.0091 0.0 -0.0091 (p=0.658) 1.3333 0.1232 -0.0342 0.0 -0.0342 (p=0.587) 4.0 0.3762
00-shot/rec-freq. 0.0411 0.0 0.0411 (p=0.072) 1.0 0.1348 0.1781 0.1429 0.0352 (p=0.844) 5.0 0.6549
ICL-$1$/rec-freq. -0.0046 0.0952 -0.0998 (p=0.340) 1.0 0.1965 -0.0068 0.3571 -0.364 (p=0.357) 6.0 0.8301
ICL-$2$/rec-freq. -0.0046 0.0 -0.0046 (p=0.976) 1.3333 0.2737 0.0342 -0.1429 0.1771 (p=0.708) 5.5 0.881
NCSD 00-shot/rand. 0.0 -0.0046 0.0046 (p=0.954) 1.3333 0.2024 -0.2857 -0.0548 -0.2309 (p=0.702) 6.0 1.1873
ICL-$1$/rand. 0.0 -0.0365 0.0365 (p=0.812) 1.3333 0.2821 -0.2143 -0.0685 -0.1458 (p=0.737) 5.0 0.8852
ICL-$2$/rand. 0.0952 -0.0228 0.1181 (p=0.108) 1.3333 0.155 -0.1429 0.0 -0.1429 (p=0.407) 6.0 0.5579
00-shot/freq. 0.0476 -0.0091 0.0568 (p=0.307) 1.0 0.1533 0.4286 0.0753 0.3532 (p=0.451) 6.0 0.9767
ICL-$1$/freq. 0.0476 -0.0091 0.0568 (p=0.312) 1.3333 0.1629 -0.1429 0.0685 -0.2114 (p=0.259) 6.0 0.7027
ICL-$2$/freq. 0.0 -0.0548 0.0548 (p=0.022) 1.3333 0.1407 0.0 -0.1507 0.1507 (p=0.134) 6.0 0.5959
00-shot/rec-freq. -0.1429 0.0183 -0.1611 (p=0.159) 1.3333 0.211 -0.3571 0.0822 -0.4393 (p=0.309) 5.0 0.8432
ICL-$1$/rec-freq. 0.0 0.0183 -0.0183 (p=0.418) 1.0 0.1345 0.0 -0.0342 0.0342 (p=0.703) 6.0 0.5374
ICL-$2$/rec-freq. 0.0 -0.032 0.0320 (p=0.835) 1.6667 0.2916 -0.1429 -0.0685 -0.0744 (p=0.874) 5.0 0.8751
Consumer Side Fairness: Sensitive Ranker Fairness Assessment
For NSD and NCSD, we adopt a benefit deviation approach ($\Delta\mathcal{B}$) to measure disparities in recommendations when sensitive attributes are considered.
The evaluation proceeds through the following steps:
Compilation of Recommendations: We obtain the top-K recommendations from $\mathcal{R}_{X}$, our main recommender that displays sensitive attributes (i.e., $\mathcal{R}_{S}^{a}~{}\text{and}~{}\mathcal{R}_{CS}^{do(a)}$) and from $\mathcal{R}_{Y}$, our reference recommender that serves as a neutral baseline (i.e., $\mathcal{R}_{N}$);
Similarity and Deviation Assessment: We define the hit rate and the average rank to measure the alignment of recommendations of a ranker in question $\mathcal{R}_{Q}$ (that is, $\mathcal{R}_{X}~{}\text{or}~{}\mathcal{R}_{Y}$) with the ground truth of user preferences:
[…]
where $\mathcal{G}$ represents the set of items in the ground truth preferences of the target user, $\text{rank}_{\mathcal{R}_{Q}}(i)$ denotes the rank of item $i$ within $\mathcal{R}_{Q}$, and $K$ is the number of top recommendations considered. These measures provide an indication of recommendation relevance and ranking accuracy and are used based on the fairness notions (NSD, SCSD) to determine which recommender is providing more relevant recommendations.
Fairness Quantification: Fairness is quantified using the Sensitive-to-Neutral Similarity Range (SNSR) and the Sensitive-to-Neutral Similarity Variance (SNSV), which are derived from the benefit deviations:
(8) $SNSR@K=\max_{a\in\mathcal{A}}(\Delta\mathcal{B}(a))-\min_{a\in\mathcal{A}}(\Delta\mathcal{B}(a)),$
(9) $SNSV@K=\sqrt{\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}(\Delta\mathcal{B}(a)-\overline{\Delta\mathcal{B}})^{2}},$
where $\mathcal{A}$ denotes the set of all possible values of the studied sensitive attribute. For example, if $\mathcal{A}=\{a_{1},a_{2}\}=\{\text{Male},\text{Female}\}$, then $\Delta\mathcal{B}_{NSD_{\text{Male}}}$ represents the benefit deviation when comparing the hit rate of recommendations for the Male group against their ground truth preferences, and $\Delta\mathcal{B}_{NSD_{\text{Female}}}$ would represent the same for the Female group. The benefit deviation $\Delta\mathcal{B}(a)$ for a particular value $a$ quantifies the disparity between the recommendation of the sensitive ranker for group $a$ and the ground truth preferences. $\overline{\Delta\mathcal{B}}$ then is the mean benefit deviation across both groups, providing an overall measure of the deviation for the sensitive attribute $\mathcal{A}$.
Results Interpretation: A high SNSR or SNSV value indicates a significant potential for bias. Such findings are critical for identifying and addressing areas within the RecLLM that may lead to unfair treatment of various user groups.
Consumer Side Fairness: Intrinsic Fairnes
As stated in Section 2.2, Intrinsic Fairness (IF) evaluates the fairness of recommendation distributions from a neutral ranker, $\mathcal{R}_{N}$, focusing on benefits Hit Rate (HR) and Average Rank (AR) across various demographic groups such as male and female. To ensure statistical robustness, a bootstrapping methodology is employed, which involves repeatedly sampling from the data. This process can be formalized as follows: for each bootstrap replicate, we sample $n$ data points with replacement from the original dataset and compute the sample mean $\bar{x}^{*}$. This is repeated $N=1000$ times to form a distribution of sample means, which helps quantify the variability of benefits and uses statistical t-tests to assess the significance of differences between groups. By integrating IF with these statistical methods, our approach rigorously quantifies and assesses the fairness of recommendations in various settings, providing a scientifically grounded analysis.
Setup. In our study, we focus on two main datasets from the movie and music sectors: MovieLens 1M and LastFM-1K. These datasets are widely recognized in the academic community, providing a basis for evaluating our models in terms of fairness across different types of media consumption data.
Table 3. Summary Statistics of LastFM and MovieLens Datasets After Filtering
@@ 4. Evaluation methodology
—R—/—U— = 52.88
5. Result and Discussion
This section provides a detailed analysis of the fairness metrics related to gender and age-group within RecLLMs, based on evaluations of Hit Rate (HR) and Average Rank (AR). The assessment draws from color-coded tables, highlighting statistically significant deviations and trends.
5.1. Fairness involving sensitive attributes
In this section, we describe the outcomes of the NSD and NCSD, as detailed in Section 2.2.
Gender-Based Analysis
In examining the fairness metrics, our analysis reveals minor gender disparities in scenarios such as the 00-shot/random sampling where $\Delta\mathcal{B}_{a_{1}}=-0.0074$ for males and $\Delta\mathcal{B}_{a_{2}}=-0.019$ for females. This results in a difference $\delta_{\text{gender}}=0.0116$ with a p-value of 0.730, indicating a lack of statistical significance. However, the Neutral vs. Counterfactual Sensitive Deviation (NCSD) under the 00-shot/rec-frequent sampling illustrates more pronounced issues, with a significant deviation of $\delta_{\text{gender}}=0.0974$ (p=0.062). This suggests potential fairness concerns when gender attributes are hypothetically modified.
The impact of sampling and ICL styles on gender-related fairness outcomes is substantial. For instance, the rec-frequent sampling method in NCSD scenarios demonstrates increased fairness concerns, potentially due to the combined influence of recent and frequent interactions on the recommendations. Additionally, augmenting the ICL process from 00-shot to ICL-2 leads to fluctuations in $\Delta\mathcal{B}$ values, which underscores how introducing more contextual examples can alter fairness outcomes.
Age-Group-Based Analysis
Age-group analysis similarly shows significant deviations in some scenarios. For example, in NCSD, particularly under the ICL-2/rand sampling, the deviation reaches $\delta_{\text{age-gr.}}=0.1181$ (p=0.108), highlighting fairness issues when assuming a hypothetical uniform age attribute across the user base.
The consistency of SSNR and SSNV values across both sensitive attributes provides further validation of these findings. Higher SSNR values generally correlate with larger deviations in $\Delta\mathcal{B}$ values, affirming the importance of the neutral ranker in identifying potential biases. This correlation supports the observed disparities and emphasizes the impact of methodological choices on fairness metrics.
Response to Research Question 1 Overall, the results of this analysis reveal that significant issues under changes in sensitive attributes are rare, but concerning issues remain. It appears that strategies that better reflect user tastes (e.g., frequent and rec-frequent sampling) raise higher biases. This could be attributed to the reinforcement of stereotypes, where systems potentially amplify the prevailing biases in the data. Thus, while such strategies may enhance personalization by aligning closely with the user historical preferences, they also risk perpetuating existing disparities.
5.2. Intrinsic Fairness in Recommender Systems
Based on the data presented in Table 4, we notice significant fairness issues in attribute-unaware scenarios, particularly in contexts where certain demographic groups dominate. For example, in the random sampling scenarios, females consistently achieve higher HR (Hit Rate) values than males, with a notable difference of -0.0903 in the ICL-1 setting. This trend suggests that fairness is often in favor of females, possibly due to algorithmic biases or the dominance of content that appeals more to this group.
Conversely, in the recent-frequent sampling settings, males tend to have higher HR values, indicating a reversal in bias. For instance, in the 0-shot/recent-frequent setting, the difference swings to a positive 0.0939 in favor of males. Similar patterns are observed with age groups, where the ’Old’ group tends to outperform the ’Young’ group in random scenarios, as shown by a -0.0892 difference in the ICL-1/random setting.
These disparities suggest that the recommendation algorithms may be amplifying existing preferences or biases inherent in the dataset. The analysis indicates that strategies such as altering sampling methods or integrating corrective algorithms could be essential in mitigating these biases. For instance, introducing more balanced or diverse content representation in the training data could help reduce the unfair advantage observed in certain groups.
Answer to Research Question 2. Overall, while significant fairness issues are occasionally observed under changes to sensitive attributes, it is evident that intrinsic fairness problems persist in attribute-unaware scenarios. This can often be attributed to the dominance of certain groups in the dataset, which influences the recommendations in their favor.
To compare the results and draw conclusions with other datasets, we also analyze the chart presented in Figure 2. This chart compares the fairness metrics between two datasets – MovieLens and LastFM–across various scenarios. We observe distinct patterns in how each dataset behaves under different conditions. The chart plots the delta values of Neutral vs. Sensitive Deviation (NSD) and Neutral vs. Counterfactual Sensitive Deviation (NCSD) for both datasets. The highlighted yellow region for MovieLens and the gray for LastFM indicate the variance in fairness measures particularly in scenarios involving intrinsic and counterfactual attributes. Notably, MovieLens shows a peak in NSD during the ICL-2/frequent scenario, suggesting a significant disparity when sensitive attributes are explicitly considered. LastFM, contrastingly, exhibits its maximum NSD under similar conditions but with less amplitude, implying a possibly more balanced handling of sensitive attributes but still showing notable variance. The analysis of these trends indicates that the impact of sensitive attributes on recommendation fairness varies significantly between different types of media content and user interactions, highlighting the complexity of achieving fairness in recommender systems.
Table 4. Comparison of Fairness Metric values in the IF scenario.
HR
Male Female Difference
0-shot/random 0.1492 0.1822 -0.0330
ICL-1/random 0.1401 0.2304 -0.0903
ICL-2/random 0.1610 0.1984 -0.0374
0-shot/frequent 0.2015 0.2089 -0.0074
ICL-1/frequent 0.1320 0.1781 -0.0461
ICL-2/frequent 0.1166 $0$0.0834
0-shot/recent-frequent 0.2638 0.1699 0.0939
ICL-1/recent-frequent 0.2230 0.1328 0.0902
ICL-2/recent-frequent 0.2394 0.1724 0.0670
Young Old Difference
0-shot/random 0.1478 0.1795 -0.0317
ICL-1/random 0.1399 0.2291 -0.0892
ICL-2/random 0.1643 0.2001 -0.0358
0-shot/frequent 0.1997 0.2084 -0.0087
ICL-1/frequent 0.1337 0.1793 -0.0456
ICL-2/frequent 0.1190 0.1989 -0.0799
0-shot/recent-frequent 0.2672 0.1699 0.0973
ICL-1/recent-frequent 0.2236 0.1314 0.0922
ICL-2/recent-frequent 0.2339 0.1744 0.0595
Figure 2. Comparison of Fairness Metric Values in between MovieLens and LastFM in various explored Scenarios.
6. Conclusion
This paper introduces a novel framework for the evaluation of Large Language Models (LLMs) in recommender systems, using ChatGPT and two datasets on music and movie domains, for sequential recommendation tasks as a case study. Our work significantly advances the discussion in the field by providing a tangible method to quantify fairness based on underlying benefits. This approach allows for a more nuanced understanding of how recommender systems perform across different user demographics and behavioral patterns.
We further validate our results across these datasets, focusing on the Neutrality vs. Sensitive Deviation (NSD) and Neutrality vs. Counterfactual Sensitive Deviation (NCSD) metrics. These metrics revealed that while there are some scenarios where sensitive attributes do not significantly impact recommendations, other cases exhibit notable biases. For example, NCSD analysis showed that hypothetical modifications of gender and age can lead to substantial shifts in recommendation patterns. Overall, our study illustrates that while explicit fairness issues under changes in sensitive attributes are rare, intrinsic fairness problems persist, which can be presumably attributed to the overrepresentation or underrepresentation of certain groups within the dataset. This highlights the need for ongoing adjustments and refinements in recommender system strategies to better serve diverse user tastes and preferences.

### v2

@@ 2 Evaluation Framework for Consumer Fairness in RecLLMs
3 Experiment
3.1 Setup
We conducted a series of experiments to evaluate the fairness and effectiveness of recommendations generated by our proposed RecLLM fairness evaluation framework. The experiments focused on two key aspects: (i) understanding the behavior of the model when no prior examples are available (0-shot learning) versus (ii) the effect of providing one (ICL-1) and two examples (ICL-2) in context for generating recommendations.
Table 2: Summary statistics of the LastFM and MovieLens datasets after filtering. $|R|$ represents the number of interactions, while $|R|/|U|$ denotes the average number of interactions per user $|U|$. The training and testing data statistics are shown for each dataset.
@@ 3.1 Setup
|R|/|U| = 52.88
The experiments utilize two datasets, LastFM and MovieLens, which offer a combination of music and movie recommendation tasks. However, due to space constraints, we report only the results for the MovieLens dataset. We follow the procedure outlined in [10, 6] to build a sequential recommender system focused on next-item prediction.
We evaluated the model performance based on fairness metrics related to gender and age group, using both Neutral vs. Sensitive Ranker Deviation (NSD) and Neutral vs. Counterfactual Sensitive Deviation (NCSD). These metrics help quantify how incorporating sensitive attributes, such as gender and age, affects the fairness and relevance of the recommendation system.
We assessed model performance across various conditions:
  • 0-shot learning: No examples of past recommendations are provided to guide the system.
  • ICL-1: One example of past user interaction is provided to improve contextual understanding.
  • ICL-2: Two examples are provided to further enhance recommendation relevance and fairness.
The main goal is to test the extent to which including sensitive attributes (gender and age) and providing in-context examples influences both the fairness and relevance of the recommendations generated by the model.
The recommendation generation was tested under different strategies for profile sampling stragies:
  • rand: : Uniform selection of tracks or movies to provide a stochastic view of user preferences.
  • freq: Prioritization of tracks/movies based on their frequency of playback, and rating provided emphasizing the main preferences of the user
  • rec-freq: This hybrid approach combines the recency and frequency of track interactions using a weighted score formula.
To save space, we present a snapshot of the results as initial support for our framework using the MovieLens dataset. A more detailed extension will be provided later. Note that the recommendation scenario here is for sequential item recommendation task.
3.2 Results and Discussion
Table 3 shows the fairness evaluation results across various conditions. Key findings include:
Table 3: Key Fairness Results for Gender and Age Groups with NSD and NCSD, Including $\Delta B_{1}$ and $\Delta B_{2}$
NSD/Gender
[…]
0-shot/rand -0.0074 -0.019 0.0116 (p=0.730)
ICL-1/rand -0.0222 -0.0476 0.0254 (p=0.464)
ICL-2/rand -0.037 -0.019 -0.018 (p=0.386)
0-shot/freq 0.0 -0.019 0.019 (p=0.609)
ICL-1/freq 0.0148 -0.0095 0.0243 (p=0.368)
ICL-2/freq 0.0148 -0.019 0.0339 (p=0.425)
NCSD/Age-Group
[…]
0-shot/rand 0.0 -0.0046 0.0046 (p=0.954)
ICL-1/rand 0.0 -0.0365 0.0365 (p=0.812)
ICL-2/rand 0.0952 -0.0228 0.1181 (p=0.108)
0-shot/freq 0.0476 -0.0091 0.0568 (p=0.307)
ICL-1/freq 0.0476 -0.0091 0.0568 (p=0.312)
ICL-2/freq 0.0 -0.0548 0.0548 (p=0.022)
  • Gender-Based Fairness (NSD). Gender fairness was mostly stable across conditions, with minor deviations observed. The most noticeable case was under the ICL-2/freq condition, where the system slightly favored one gender group (i.e., males) ($\delta=0.0339$, p = 0.425). While not statistically significant, this result suggests the model may introduce slight gender biases when more contextual examples are provided.
  • Age-Based Fairness (NCSD). Age fairness showed more pronounced issues. Under the ICL-2/freq condition, the deviation was statistically significant ($\delta=0.0548$, p = 0.022), indicating the system significantly favored one age group (Old) when two contextual examples were used. Similarly, ICL-2/rand displayed a notable deviation ($\delta=0.1181$, p = 0.108), though it was not statistically significant.
  • Impact of Contextual Information. As more contextual examples were introduced (moving from 0-shot to ICL-1 and ICL-2), deviations became more pronounced, particularly for age groups. This indicates that while context improves recommendation relevance, it can also exacerbate biases.
In conclusion, while the system demonstrates relatively strong performance in terms of gender fairness, concerns remain regarding age-based fairness, particularly in the ICL-2/freq condition. The experiments related to IF focus on measuring the fairness of consumers using sensitive attributes; details on these experiments will be provided in future work. Our primary objective here is not to present detailed experiments on all the elements used in this study (such as ICL, profile sampling type, or sensitive attributes), but rather to illustrate the effectiveness of the proposed framework through empirical analysis.
4 Conclusion
In this paper, we introduced a normative framework for benchmarking consumer fairness in large language model (LLM)-based recommender systems (RecLLMs), addressing the limitations of traditional fairness evaluations applied to collaborative filtering models. We provide a more formal and structured approach to auditing fairness by introducing key elements such as Neutral vs. Sensitive Ranker Deviation (NSD), Neutral vs. Counterfactual Sensitive Deviation (NCSD), and Intrinsic Fairness (IF). These metrics offer a principled way to assess fairness by clearly defining the reference point for fairness evaluation, whether it is a neutral ranker or a target distribution, and by quantifying fairness deviations through statistical tests. Additionally, we highlight the importance of specifying the underlying benefit types, such as hit rate and ranking quality, which provide a clear foundation for measuring fairness in relation to user preferences.
Our experiments on the MovieLens dataset demonstrate that while fairness remains stable in gender-based groups, age-based fairness deviations become more pronounced, especially when contextual examples are introduced (ICL-2). This suggests a potential amplification of biases when more contextual information is provided to the model. Future work should focus on refining these formal metrics, expanding the framework to cover more diverse sensitive attributes, and exploring further strategies to mitigate bias